An asset with initial cost base of 6 million was bought 5 years ago.At the end of the current year (fifth year) the asset is sold by 1.1 million. The asset belongs to the 7 years MACRS property class. Knowing that the gain (loss) tax rate 40%,what is the net proceeds resulted from selling the asset?


1
Expert's answer
2021-04-29T07:35:35-0400

the initial cost of the fixed asset has been written off: 6million

the amount of depreciation on the realized fixed asset is written off:

"6\times\frac{1}{7}=0.857"

"0.857\times5=4,29"

the asset is sold by 1.1 million

net proceeds=1.1+4.29-6-(1.1+4.29-6)\times0.4=-0.37